OpenStudy (lwaluko):

HELP PLease!! Lisa's house is represented on the graph with the point (3,2). Tina's house is 12 miles north and 5 miles east of Lisa's house and is represented with the point (8,14). The shortest distance from Lisa's house to Tina's house is ________ miles.

OpenStudy (sunnnystrong):



OpenStudy (sunnnystrong):

Recall that: \[a^2+b^2=c^2\] @lwaluko ... what do you think the shortest distance is?

OpenStudy (lwaluko):

17?

OpenStudy (sunnnystrong):



OpenStudy (sunnnystrong):

\[(12)^2+(5)^2=c^2\] Solve for c --> \[c=\sqrt{169}\] c=13miles. So.. she can travel up 12 miles and to the right 5 miles or just travel 13 miles

OpenStudy (lwaluko):

Ohh

OpenStudy (lwaluko):

Thanks so much!